Your brother leaves on a trip, forgetting his suitcase. You know that he normally drives at a speed of 55mph. You do not discover the suitcase until 1 hour after he has left. If you follow him at a speed of 65mph, how long will it take you to catch up with him?

Answer by stanbon(75887)

Your brother leaves on a trip, forgetting his suitcase. You know that he normally drives at a speed of 55mph.
You do not discover the suitcase until 1 hour after he has left. If you follow him at a speed of 65mph, how long will it take you to catch up with him?
----------------------------------------
Brother DATA:
rate = 55 mph; time = x+1 ; distance = 55(x+1) miles
-----------------------------
Your DATA:
rate = 65 mph ; time = x hrs: distance = 65x miles
------------------------
Equation:
55(x+1) = 65x
55x+55 = 65x
10x = 55
x = 5.5 hrs (time required to catch your brother)
==================
Cheers,
Stan H.